# SynParaSpeech: Automated Synthesis of Paralinguistic Datasets for Speech Generation and Understanding

### v1

@@ 2.1 SynParaSpeech Labeled Text Synthesis
To add paralinguistic tags, we used a mainstream large language model (LLM) Deepseek Chat V3 [18] to build a tagged text dataset. We input the aforementioned transcriptions into the LLM, prompting it to select the optimal tag from [laugh], [sigh], [gasp], [throat clearing], [pause], [tsk] without altering the original text. The LLM also found optimal insertion spots in each VAD-split clip (e.g., sentence start/clip end). Finally, we got text data with paralinguistic labels.
@@ 4.1.1 Paralinguistic TTS
In the paralinguistic TTS experiments, we employed SynParaSpeech to conduct supervised fine-tuning (SFT) on state-of-the-art open-source TTS models, namely the autoregressive CosyVoice2 [1] and the non-autoregressive F5-TTS [21]. For comparison, we also considered the NVS dataset [12], a large and diverse paralinguistic resource collected from real-world media and automatically annotated. To ensure comparability, we selected four common paralinguistic labels—[laugh], [sigh], [gasp], and [throat clearing]—that are supported by both SynParaSpeech and NVS. We kept 2% of the data randomly as a validation set, the rest for training. For evaluation, the LLM DeepSeek V3 [18] was used to rewrite seed-tts-eval texts (with paralinguistic labels inserted), making 100 inference texts per category.
@@
$\Delta(\mu_{\mathrm{p}},\mu_{\mathrm{o}};y)$ $=\log\frac{\pi_{\mathrm{ref}}(\mu_{\mathrm{p}}\mid y)}{\pi_{\theta}(\mu_{\mathrm{p}}\mid y)}-\log\frac{\pi_{\mathrm{ref}}(\mu_{\mathrm{o}}\mid y)}{\pi_{\theta}(\mu_{\mathrm{o}}\mid y)},$ (1)
$\mathcal{L}_{\text{dpo}}$ $=-\mathbb{E}_{y}\left[\log\sigma\left(\beta\cdot\Delta(\mu_{\mathrm{p}},\mu_{\mathrm{o}};y)\right)\right],$
@@
For training, CosyVoice2’s language model is optimized with Adam at a learning rate of $1\times 10^{-5}$ for 50 epochs, with early stopping (patience 10), a constant learning rate, and a 2500-step linear warm-up. Training stability is maintained via gradient clipping (threshold 5), gradient accumulation over 2 steps, and dynamic batching with a maximum of 2000 frames. Following the NVS setup, F5-TTS is trained for 400 epochs with a learning rate of $1\times 10^{-4}$, a cosine annealing scheduler with 1000 warm-up updates, and a frame-based batch size of 30,000 per GPU. To initialize new paralinguistic tokens, we use embeddings from the RoBERTa-based [23] text encoder of CLAP [24], aligning dimensions via interpolation.
Model performance was assessed using both objective and subjective metrics. Objective measures include character error rate (CER) for intelligibility, UTMOSv2 [25] for speech quality, and SECS for timbre similarity. Subjective evaluation employed 5-point ratings: SMOS for speaker similarity, NMOS for naturalness, QMOS for overall audio quality, and PMOS for the paralinguistic quality. For MOS ratings, 21 volunteers participated in a double-blind evaluation.
@@ 4.1.2 Paralinguistic Event Detection
In the paralinguistic event detection experiments, we applied SynParaSpeech-based prompt tuning to multimodal large language models (MLLMs) and analyzed the effect of context size. Following MMSU [26], Kimi Audio [27] was chosen for its strong paralinguistic reasoning, and Qwen 2.5 Omni [28] for its superior paralinguistic perception. Both models were evaluated with and without SynParaSpeech prompt tuning. For each paralinguistic category, 100 audio–text pairs were used for testing, while 1, 3, 5, and 7 pairs were sampled as context prompts. Performance was measured using accuracy, macro F1 score, and CER.
@@ 4.2.2 Paralinguistic Event Detection Experimental Results
Table 4 highlights the effect of SynParaSpeech prompt tuning and context quantity on paralinguistic event detection. First, the results confirm the utility of SynParaSpeech: for both Kimi Audio and Qwen 2.5 Omni, introducing SynParaSpeech prompts substantially improves performance compared with the no-context baseline, especially in terms of accuracy and macro F1 score. This verifies that the SynParaSpeech dataset provides effective supervision signals for paralinguistic perception and reasoning.
Second, the number of contextual examples exerts a critical influence. Performance generally rises with the increase in contextual quantity, and peak performance gains are observed under the 5-shot setting for both models. However, excessive context (e.g., 7-shot) fails to bring further improvements and may even lead to performance degradation. This indicates a trade-off between providing informative prompts and avoiding model overload caused by redundant input.
Overall, these findings demonstrate that SynParaSpeech enhances the model’s capability in paralinguistic event detection, while an optimal quantity of contextual examples is essential to fully leverage its advantages.

### v3

@@ 2.1 SynParaSpeech Labeled Text Synthesis
To incorporate paralinguistic tags, we used a mainstream large language model (LLM) Deepseek V3 [18] to build a tagged text dataset. We input the transcriptions into the LLM, prompting it to select the most appropriate tag from [laugh], [sigh], [gasp], [throat clearing], [pause], [tsk] without altering the original text and insert it at the boundary of relevant text segments. Ultimately, this process yielded a text dataset annotated with paralinguistic tags.
@@ 4.1.1 Paralinguistic TTS
In the paralinguistic TTS experiments, we employed SynParaSpeech to conduct supervised fine-tuning (SFT) on state-of-the-art open-source TTS models, namely the autoregressive CosyVoice2 [1] and the non-autoregressive F5-TTS [21]. For comparison, we also considered the NVS dataset [12], a large and diverse paralinguistic resource collected from real-world media and automatically annotated. To ensure comparability, we selected four common paralinguistic labels—[laugh], [sigh], [gasp], and [throat clearing]—that are supported by both SynParaSpeech and NVS. We kept 2% of the data randomly as a validation set, the rest for training. For evaluation, the LLM DeepSeek V3 [18] was used to rewrite Seed-TTS-Eval [22] texts (with paralinguistic labels inserted), making 100 inference texts per category.
@@
$\mathcal{L}_{\text{dpo}}=-\mathbb{E}_{y}\left[\log\sigma\left(\beta\cdot\log\frac{\pi_{\theta}(\mu_{p}\mid y)\cdot\pi_{\text{ref}}(\mu_{o}\mid y)}{\pi_{\theta}(\mu_{o}\mid y)\cdot\pi_{\text{ref}}(\mu_{p}\mid y)}\right)\right]$ (1)
@@
For training, CosyVoice2’s language model is optimized with Adam [24] at a learning rate of $1\times 10^{-5}$ for a maximum of 50 epochs, with early stopping (patience 10), a constant learning rate, and a 2500-step linear warm-up. Training stability is maintained via gradient clipping (threshold 5), gradient accumulation over 2 steps, and dynamic batching with a maximum of 2000 frames. Following the NVS setup, F5-TTS is trained for 400 epochs with a learning rate of $1\times 10^{-4}$, a cosine annealing scheduler with 1000 warm-up updates, and a frame-based batch size of 30,000 per GPU. To initialize new paralinguistic tokens, we use embeddings from the RoBERTa-based [25] text encoder of CLAP [26], aligning dimensions via interpolation.
Model performance was assessed using both objective and subjective metrics. Objective measures include character error rate (CER) for intelligibility, SECS for speaker similarity, and UTMOSv2 [27] for speech quality. Subjective evaluation employed 5-point ratings: PMOS for the paralinguistic quality, NMOS for naturalness, SMOS for speaker similarity, and QMOS for overall audio quality. For MOS ratings, 21 volunteers participated in a double-blind evaluation.
@@ 4.1.2 Paralinguistic Event Detection
In the paralinguistic event detection experiments, we applied SynParaSpeech-based prompt tuning to multimodal large language models (MLLMs) and analyzed the effect of context size. Following MMSU [28], Kimi Audio [29] was chosen for its strong paralinguistic reasoning, and Qwen 2.5 Omni [30] for its superior paralinguistic perception. Both models were evaluated with and without SynParaSpeech prompt tuning. Each paralinguistic category used 100 test pairs with varying context prompts. Performance was measured using accuracy, macro F1 score, and CER.
@@ 4.2.2 Paralinguistic Event Detection Experimental Results
Table 4 shows the impact of SynParaSpeech prompt tuning and context quantity on paralinguistic event detection. Notably, due to Kimi Audio’s poor support for long-context input versus Qwen 2.5 Omni’s strong support, the context quantity for Kimi Audio refers to the total number of prompts, while that for Qwen 2.5 Omni corresponds to the number per paralinguistic category. SynParaSpeech consistently improves accuracy and macro F1 over the no-context baseline for both models, confirming its value for paralinguistic perception and reasoning. Context quantity also matters: performance improves with more examples, peaking at 5-shot; beyond this (e.g., 7-shot), the improvements stop or even get worse. This reflects a trade-off between informative prompts and avoiding input overload. Overall, SynParaSpeech enhances paralinguistic event detection capability, and an optimal context size is essential to maximize its benefits.
@@ 5 Conclusion
6 Acknowledgement
The work was supported by the National Natural Science Foundation of China (NSFC) (No. 62271083), the Key Project of the National Language Commission (No. ZDI145-81), and the Fundamental Research Funds for the Central Universities (No. 2023RC73).